### v1

# TinyTim: A Family of Language Models for Divergent Generation

Abstract
This work introduces TinyTim, a family of large language models fine-tuned on James Joyce’s ‘Finnegans Wake’. Through quantitative evaluation against baseline models, we demonstrate that TinyTim V1 produces a statistically distinct generative profile characterized by high lexical diversity and low semantic coherence. These findings are interpreted through theories of creativity and complex problem-solving, arguing that such specialized models can function as divergent knowledge sources within more extensive creative architectures, powering automated discovery mechanisms in diverse settings.
1 Introduction
Large Language Models (LLMs) based on the Transformer architecture (vaswani2017attention, ) have demonstrated powerful capabilities in synthesizing statistically likely patterns from vast data (brown2020language, ). However, this strength is also a fundamental limitation, constraining them to convergent, ‘mean-reverting’ outputs that inhibit the generation of genuinely novel hypotheses (dweck2023creativity, ). This issue reflects long-standing critiques of artificial reason, where formal systems struggle with the ambiguity and creative extrapolation inherent in human cognition (dreyfus1992, ; piantadosi2022meaning, ).
Human creativity is often described by dual-process theories that posit an interplay between focused, convergent thinking for evaluation and associative, divergent thinking for ideation (guilford1967nature, ; sternberg1985implicit, ). This cognitive duality is linked neuroscientifically to the dynamic coupling of the brain’s Executive Control Network (ECN) and Default Mode Network (DMN) (beaty2016creative, ; andrews201813, ; chen2025dynamic, ; LUCHINI2023195, ). While standard LLMs excel at tasks analogous to convergent thinking, they lack an intrinsic mechanism for the spontaneous, associative thought characteristic of divergence. To address this deficit, we engineered a divergent generative process into a language model by fine-tuning it exclusively on James Joyce’s ‘Finnegans Wake’—a text that is itself a product of extreme associative thought.
2 Methodology
2.1 Data and Training
We fine-tuned the ‘TinyLlama-1.1B-Chat-v1.0’ model on the complete text of ‘Finnegans Wake’ (approx. 1.5MB). The text was preprocessed into 100-word segments to preserve its associative structure. The model, named TinyTim, was trained using a standard causal language modeling objective. TinyTim V1 has been available on HuggingFace for over a year 111hf.co/npc-worldwide/TinyTimV1, and has been downloaded more than 750 times in that time span.
In this work we showcase and characterize the V1 of the TinyTim family, and we plan to soon train and release an instruction-tuned version that can not only generate text like Joyce but can answer questions and be helpful as well to further facilitate creative processes.
2.2 Evaluation Framework
To quantify TinyTim’s generative profile, we compared it against a suite of baseline models trained for coherent responses: ‘qwen3:0.6b’, ‘llama3.2’, and ‘gpt-5-mini’. We generated responses from each model using a set of 10 creative prompts. Each response was evaluated using syntactic and semantic metrics:
  • Syntactic Metrics: Unique Word Ratio, Average Word Length, Token Diversity (Shannon entropy), and Sentence Complexity.
  • Semantic & Content Metrics: Semantic Similarity to the prompt (via sentence embeddings), Readability (Flesch-Kincaid Grade Level), and Sentiment (VADER compound score).
3 Results
The generative profiles of TinyTim and the baseline models were evaluated across a suite of syntactic, lexical, and semantic metrics. The analysis reveals a statistically significant and functionally distinct separation between TinyTim and all conventionally trained models.
From an initial set of 2400 generated samples, 1013 were deemed valid for analysis after filtering out malformed or single-word outputs. The final sample sizes were: TinyTim (n=714), ‘gpt-5-mini’ (n=99), ‘llama3.2’ (n=100), and ‘qwen3:0.6b’ (n=1100). A Kruskal-Wallis test confirmed a significant difference ($p<.0001$) among the model groups for all metrics. Subsequent pairwise Mann-Whitney U tests with Bonferroni correction found that TinyTim was statistically different from all three baseline models on every metric.
Analysis of lexical sophistication metrics reveals a fundamental difference in how the models use vocabulary. As shown in Figure 1, while ‘gpt-5-mini’ possesses the largest raw vocabulary size, TinyTim demonstrates a radical degree of lexical invention.
Its Hapax Legomena Ratio, the proportion of words used only once, is over 50% higher than ‘gpt-5-mini’ (0.643 vs. 0.413), and its Yule’s K score—a robust measure of lexical richness that accounts for repetition—is over four times greater (208 vs. 47). This indicates TinyTim’s generative process is not based on retrieving a wide range of common words from a large vocabulary, but on constantly producing novel, rare, and often neologistic terms from a more constrained set of inputs. The baseline models are sophisticated retrievers; TinyTim is a lexical inventor.
Figure 1: Measures of lexical sophistication. TinyTim is a clear outlier in lexical invention (Hapax Ratio, Yules K), distinguishing between creative generation and simple vocabulary size.
The distributional profiles of the models, shown in Figure 2, highlight the difference between convergent and divergent systems. The baseline models (‘gpt-5-mini’, ‘llama3.2’, ‘qwen3:0.6b’) produce outputs with tight, narrow, and predictable distributions for all metrics. This is the statistical signature of a reliable, convergent system optimized for consistency.
In contrast, TinyTim’s distributions are extremely wide, heavily skewed, and possess long tails. This is most apparent in Unique Word Ratio, where its distribution spans a range nearly double that of any baseline, and in Sentence Complexity, where its outputs range from single-word fragments to sentences over 200 words long. This high variance is not a flaw, but the primary indicator of its function as an unpredictable, divergent generator whose outputs are intentionally varied and chaotic.
Figure 2: Ridge plots showing the distribution of scores for the four primary metrics. The baseline models show tight, predictable distributions. TinyTim exhibits extreme variance and long tails, indicative of a divergent generative process. The x-axis represents the score for the corresponding metric.
The relationships between metrics further illuminate the different generative strategies. Figure 3 shows the correlations between key metrics across the full dataset. The most significant finding is in the top-left panel, which plots Token Diversity against Unique Word Ratio.
The baseline models form a tight cluster in the high-diversity, low-uniqueness space. They draw from a vast, general vocabulary (high Token Diversity) but exhibit more repetition and less novelty within any single response (lower Unique Word Ratio). TinyTim occupies the opposite corner of the plot, forming a dispersed cloud in the high-uniqueness, low-diversity space. This is confirmed by the strong negative correlation for the full dataset (r=0.807r=-0.807). This illustrates the core paradox of its creativity: each individual response is highly novel (high UWR), but the overall pool of words it uses is highly specialized and constrained to its training data (low Token Diversity). This trade-off is the defining feature of its specialized, divergent function.
Figure 3: Scatter plots showing relationships between primary metrics. The top-left panel highlights the fundamental trade-off between Token Diversity and Unique Word Ratio, clearly separating TinyTim’s divergent strategy from the convergent strategy of the baseline models.
4 Discussion
The quantitative results show that fine-tuning on experimental literature produces a model with a statistically distinct, divergent generative profile. This moves the assessment of AI creativity beyond subjective claims and toward a measurable, multi-faceted profile.
4.1 Challenging Parsimony in Creative Systems
The principle of parsimony often guides the development of scientific models. However, recent work suggests that overly simplistic models can be insufficient for complex phenomena, and that complexity can be essential for scientific progress (dubova2025, ). Standard LLMs are inherently parsimonious; they seek the most probable, and thus often simplest, path through semantic space. ‘Finnegans Wake’ is an anti-parsimonious text. By training on it, TinyTim learns a complex, non-minimal generative function. This demonstrates that for tasks requiring creative exploration, a non-parsimonious model trained on complex data may be more valuable than a simple one, as it is better equipped to navigate the high-complexity spaces where novel ideas reside (boden1998creativity, ; holland1992genetic, ). This aligns with findings that unguided, random exploration can be superior to rigid, theory-driven investigation (dubova2022against, ; musslick2025automating, ). Furthermore, when paired with a convergent model, outputs from TinyTim can be harnessed as part of an integrated system of exploration, positioning it as a powerful tool for automated discovery.
4.2 Implications for Human-AI Collaboration
The utility of TinyTim lies in its ability to generate novel linguistic fragments and non-obvious connections, fostering the conditions for serendipity (mccay2015investigating, ). This raw material requires interpretation. The user’s role shifts from querying for an end to engaging in a co-creative, human-in-the-loop process (hmcf2025, ; labelyourdata2025, ). This process is analogous to Conceptual Blending, where new meaning emerges from the integration of disparate mental spaces (fauconnier2003conceptual, ). The model provides the disparate spaces; the human—or an alternative convergent LLM [^2] The way Large Language Models interpret meaning in words appears to reproduce the contextuality effects observed in humans, see Agostino et al. (2025) on quantum semantics agostino2025quantum and Wu et al. (2024) on the relative and sequential placement of words matters to LLMs and bruza2023 to read a recent description of the process observed in cognitive experiments.—provides the interpretive act. A system that allows a user to invoke a ‘divergent agent’ makes this creative strategy an accessible part of the collaborative process [^3] Indeed, TinyTim has already played a role in powering creative work: Don’t turn on the sun by giacomo catanzaro , enhancing the system’s ability to respond to context in a useful manner (lampinen-etal-2022-language, ).
5 Conclusion
This work has demonstrated that the generative properties of a language model can be fundamentally reshaped through targeted fine-tuning on radically unconventional literature. By analyzing the resultant model, TinyTim, against a suite of baselines, we have established a set of core findings regarding the nature of artificial creativity and model specialization. Our major conclusions are the following:
  1. 1.
    Fine-tuning on experimental literature quantitatively demonstrates that a language model’s intrinsic generative bias can be shifted from a convergent to a divergent cognitive style, as evidenced by a statistically significant increase in lexical invention and output variance.
  2. 2.
    The generative profile of TinyTim establishes a formal distinction between the convergent, retrieval-based sophistication of standard LLMs and a divergent, combinatorial creativity that achieves novelty by exhaustively exploring the latent space of a focused, complex domain.
  3. 3.
    The model’s unique profile serves as a proof-of-concept for a specialized "divergent knowledge source," validating the architectural principle of using heterogeneous multi-agent systems for complex problem-solving.
  4. 4.
    The model’s function as a generator of high-variance, low-coherence semantic material implies a new paradigm for human-AI interaction, shifting from a query-response dynamic to a co-creative partnership where the user (or another convergent LLM) must perform the interpretive act.